OpenStudy (anonymous):

Mira has $760 in her bank account and withdraws $40 each month. Paulina has $210 in her bank account and deposits $70 each month. In how many months will Mira and Paulina have equal amounts of money in their accounts? A. 5 B. 7 C. 8 D. 10

OpenStudy (gorv):

so mira's balance =760-40*x x= no of month 40x= total amount withdrawl in x month

OpenStudy (gorv):

paulina's balance = 210+70x x= no of month 70x=total deposit in x month

OpenStudy (gorv):

for equal balance 760-40*x=210+70x now solve it

OpenStudy (anonymous):

I want to say that the answer is B